# Collaborative Multi-Agent Optimization for Personalized Memory System

### 13 Mar 2026 v1

@@ Appendix A Algorithm
9:  $r_{\text{ans}}\leftarrow\mathbb{I}(p^{*}=p)$ $\triangleright$ Global reward (accuracy)
@@ B.1 Dataset
In the PersonaMem benchmark, each interaction history consists of 10, 20, or 60 multi-turn sessions (corresponding to 32k, 128k, or 1M tokens) to form varying context lengths. Each session contains 15–30 conversation turns, designed to naturally reveal or update user preferences. In addition, each history is paired with a suite of subsequent queries across seven types, enabling comprehensive evaluation of personalized response capabilities. For PersonaMem’s three context length settings (32k, 128k, and 1M tokens), we partition their training and validation sets based on these queries. The statistical breakdown of training and validation queries is summarized in Table 6.
To process the training set, we first identify the specific session within the full conversational history that contains explicit evidence relevant to each query $q$, and we define this session as the query’s targeted conversational history $\mathcal{H}$. We then construct $(\mathcal{H},q)$ data pairs required for generating individual MDP trajectories as the core training data. The position of each target session within the complete interaction history is specified in the original benchmark. Ground-truth evidence $\mathcal{E}_{q}$ for each query $q$ is extracted using the Seed-1.6 model and validated through human evaluation. For evaluation, multiple distinct queries may share the same full conversational history that includes multiple sessions. This design enables further evaluation in complex inference scenarios where memory construction agents operate on a full conversational history with multiple sessions, while retrieval agents are activated for subsequent queries at varying frequencies.
@@ B.2 Detailed Baselines
  • RAG (Lewis et al., 2020) answers queries by retrieving raw conversational history via embedding-based similarity matching. The embedding model employed is all-MiniLM-L6-v2, and ultra-long input contexts are uniformly partitioned into fixed-length segments, each consisting of 2,048 tokens.
@@
  • A-Mem (Xu et al., 2025) instantiates a memory system through multiple agents with dedicated prompts, including the memory retrieval construction, precursor memory evolution, and retrieval agents;
  • CAM (Li et al., 2025) builds a hierarchical memory system via multiple agents, where dedicated agents undertake node construction (via ego-centric disentanglement and node replication) and information aggregation (for hierarchical abstraction). We implement it on the PersonaMem benchmark.
  • MemoryBank (Zhong et al., 2024) construct a human-like memory for LLMs that uses a forgetting curve to remember and forget information over time. Similarly, we implement it on the PersonaMem benchmark.
@@
  • Mem1 (Zhou et al., 2026) proposes a reinforcement learning (RL) framework that empowers agents to conduct efficient multi-turn information retrieval and reasoning for query answering. The optimized agents prioritize direct memory retrieval from history, rather than memory construction.
  • Memory-R1 (Yan et al., 2025b) constructs a Memory Manager that learns optimal memory operations (ADD, UPDATE, DELETE) and an Answer Agent tasked with distilling relevant information from retrieved entries. Uniquely, the system is optimized via independent reinforcement learning (RL) in two stages: the Memory Manager is optimized first, followed by the Answer Agent.
@@ B.3 More Implementation Details
For a fair comparison, all the baselines are implemented with the same LLM backbones as our CoMAM, including Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct. For the implementation of variant “+independent RL”, each agent is trained by fusing its task-specific rewards and global accuracy rewards, with the remaining agents frozen. For the “single policy” variant, we adopt a shared policy to serve as the policy for all three agents during training and inference. The prompts for different agents are presented below. The temperature of LLMs is set to $1$ for training and $0.8$ for evaluation. The detailed hyperparameter settings for CoMAM are presented in Table 7.
@@ C.1 The Impact of Joint Optimization on Local Task Specialization
  • Global-local Collaboration: The global reward curve of “Ours” converges to 0.70, significantly higher than “w/ L’ (0.65) and “w/ G” (0.64), proving that joint optimization realizes global performance enhancement by fostering inter-agent collaboration.
Adaptive credit assignment mechanism quantifies the correlation between each agent’s local performance and the global outcome, ensuring that agents receive rewards not only for completing local tasks well but also for contributing to the global goal. This guides agents to optimize their local strategies in a way that benefits the entire system, thus achieving both local specialization and global collaboration.
@@ D.1 Difference with existing memory framework
    Existing frameworks primarily adopt independent optimization. For example, MIRIX (Wang & Chen, 2025) designs a meta-agent and six specialized agents for memory management, optimizing each via tailored prompts; Memagent (Yu et al., 2026) focuses on a single retrieval agent, training it with query-answer accuracy as the reward; Memory-R1 (Yan et al., 2025b) splits the system into a Memory Manager and Answer Agent, training them in separate stages—first optimizing the Manager with the Agent frozen, then freezing the Manager to optimize the Agent. This approach ignores inter-agent collaboration: memory construction is not adjusted based on retrieval needs, and the Agent can only adapt to a fixed Manager’s output, leading to “local optimal but global suboptimal” outcomes.
    CoMAM proposes end-to-end joint optimization via MDP regularization. We model the asynchronous execution of Extraction, Profile, and Retrieval Agents as a sequential MDP trajectory, where each agent’s output directly serves as the next agent’s input. All three agents are updated simultaneously during training, with inter-agent dependencies embedded into state transitions, ensuring they evolve collaboratively toward the global goal of query-answer accuracy.
@@
    Existing frameworks that optimize memory agents with RL primarily rely solely on final query-answer accuracy as the reward for all agents. For example, Mem-$\alpha$ and Mem1 (Zhou et al., 2026) use this single reward to train construction and retrieval agents respectively. Memory-R1 (Yan et al., 2025b) and GAM (Yan et al., 2025a) also adopt this design, using only answer accuracy to optimize their agents separately. This ambiguous credit assignment prevents agents from distinguishing their individual contributions, leading to unstable convergence.
@@ D.2 Difference with existing multi-agent system RL
  • Heterogeneous Policy vs. Homogeneous Policy: Some multi-agent reinforcement learning (RL) methods adopt a shared policy across all agents. For instance, CORY (Ma et al., 2024) proposes an RL framework featuring “Pioneer” and “Observer” agents that share a single policy model. These agents co-evolve through multi-agent collaboration and periodic role-swapping. In contrast, considering the heterogeneity and asynchrony of a memory system, CoMAM employs heterogeneous policies tailored to distinct agents.
  • Workflow: Sequential vs. Concurrent: Existing multi-agent system (MAS) methods often adopt a debate-style pipeline for reasoning tasks, such as math tasks. For instance, MAPoRL (Park et al., 2025) models agent interactions as concurrent debate, where multiple agents generate responses in parallel, focusing on persuasive collaboration for general reasoning. In contrast, CoMAM explicitly formulates agents’ executions as a sequential Markov Decision Process (MDP), embedding inter-agent dependencies into state transitions, ensuring that memory construction quality dynamically adapts to retrieval requirements and vice versa.
  • Collaboration Paradigm: Joint Optimization vs. Independent Alignment: Existing methods, such as Optimas (Wu et al., 2026), foster collaboration by aligning local rewards with global objectives via heuristic reward shaping for compound AI systems. However, it optimizes individual agents separately, leveraging locally-global aligned rewards rather than through end-to-end joint training. In contrast, CoMAM adopts a joint optimization paradigm: it regularizes the asynchronous execution of heterogeneous agents into a sequential MDP trajectory and updates all agents simultaneously.
  • Credit Assignment: Adaptive Assignment vs. Equal Allocation: Most existing systems use static credit assignment mechanisms without quantifying agents’ varying contributions: AT-GRPO (Zhao et al., 2026) distributes rewards equally across agents using fixed weights, ignoring task differences; CoMAS (Xue et al., 2026) adopts a zero-sum adversarial reward scheme that pits agents against each other, which is misaligned with the cooperative nature of memory systems; In contrast, CoMAM’s credit assignment is grounded in group-level ranking consistency (NDCG) between each agent’s local task rewards (e.g., coverage for extraction, rationality for profiling, precision for retrieval) and the global system reward (query-answer accuracy). This design quantifies each agent’s contribution to the global system. It dynamically adjusts weights according to agent performance, ensuring adaptive credit allocation that aligns local improvements with global performance.
@@ Appendix E Limitation
While CoMAM enables effective joint optimization for multi-agent memory systems, it is worth noting that our current implementation focuses on verifying the core logic of collaborative RL, with certain extended capabilities reserved for future exploration to better adapt to more complex scenarios.
1. Limited Agent Task Scope:
The framework currently leverages a compact agent set (Extraction, Profile, Retrieval) to focus on verifying the efficacy of collaborative RL for personalized memory tasks. To prioritize the core verification goal, we reserve extended agent capabilities (such as memory modification for consistency maintenance and redundant memory deletion) for future work.
2. Static MDP Regularization for Training:
Current MDP regularization models agent execution as a static sequential trajectory, enabling joint training of heterogeneous and asynchronous agents. Compared with the existing independent optimization training paradigm, this design can better enhance the collaborative capabilities of agents beyond task-specific capabilities. While effective in boosting inter-agent collaboration, we can further consider the dynamic deployment of agents in real-world scenarios, such as dynamic invocation frequencies and real-time memory updates. Future work will integrate these real-world dynamics into the joint training process to improve the consistency between training and deployment.
3. Lack of Explicit local-local Alignment:
The adaptive credit assignment is specifically designed to align local agent improvements with global system goals, enabling agents’ collaboration towards the global goal. Since all agents’ local subtasks are oriented towards the consistent global system goal, the alignment between their local tasks is implicitly modeled. In the future, we will explore explicitly modeling this local-local alignment for agents’ local subtasks, which will further boost inter-agent collaboration.

### 27 Apr 2026 v2

@@ Appendix A Algorithm
9:  $r_{\text{ans}}\leftarrow\mathbb{I}(p^{*}=p)$ $\triangleright$ Global reward (Downstream task accuracy)
@@ B.1 Dataset
In the PersonaMem benchmark, each interaction history consists of 10, 20, or 60 multi-turn sessions (corresponding to 32k, 128k, or 1M tokens) to form varying context lengths. Each session contains 15–30 conversation turns, designed to naturally reveal or update user preferences. In addition, each history is paired with a suite of subsequent queries across seven types, enabling comprehensive evaluation of the memory system. The question types include: (1) Recall user-shared facts, (2) Suggest new ideas, (3) Acknowledge latest user preferences, (4) Track full preference evolution, (5) Revisit reasons for preference updates, (6) Provide preference-aligned recommendations, and (7) Generalize to new scenarios. Compared to existing benchmarks [19, 32], PersonaMem’s strengths in long context history and different query type enable rigorous assessment of memory systems in conversational scenarios. For PersonaMem’s three context length settings (32k, 128k, and 1M tokens), we partition their training and validation sets based on these queries. The statistical breakdown of training and validation queries is summarized in Table 7.
In the LongMemEval benchmark [32], each sample consists of a scalable multi-session user–assistant interaction history, where information is gradually revealed and may be revised across sessions. The benchmark contains 500 manually created questions and defines two standard settings: LongMemEval-S with about 115K tokens per sample, and LongMemEval-M with 500 sessions and roughly 1.5M tokens. Its evaluation targets five core memory abilities—information extraction, multi-session reasoning, temporal reasoning, knowledge updates, and abstention—and the released question categories include single-session-user, single-session-preference, single-session-assistant, multi-session, temporal-reasoning, and knowledge-update. Compared with benchmarks that mainly test long-context recall, LongMemEval more explicitly evaluates cross-session reasoning and dynamic user-state tracking in realistic assistant settings.
To process the training set, we first identify the specific session within the full conversational history that contains explicit evidence relevant to each query $q$, and we define this session as the query’s targeted conversational history $\mathcal{H}$. We then construct $(\mathcal{H},q)$ data pairs required for generating individual MDP trajectories as the core training data. The position of each target session within the complete interaction history is specified in the original benchmark. Ground-truth evidence $\mathcal{E}_{q}$ for each query $q$ is extracted using the Seed-1.6 model and validated through human evaluation. We emphasize that $\mathcal{E}_{q}$ is only used for computing training rewards and is not available during inference, ensuring no information leakage. While obtaining $\mathcal{E}_{q}$ involves additional annotation effort, this cost is incurred only during training and does not affect inference-time efficiency, making it a practical and acceptable design choice. For evaluation, multiple distinct queries may share the same full conversational history that includes multiple sessions. This design enables further evaluation in complex inference scenarios where memory construction agents operate on a full conversational history with multiple sessions, while retrieval agents are activated for subsequent queries at varying frequencies.
@@ B.2 Detailed Baselines
  • RAG [14] answers queries by retrieving raw conversational history via embedding-based similarity matching. The embedding model employed is all-MiniLM-L6-v2, and ultra-long input contexts are uniformly partitioned into fixed-length segments, each consisting of 2,048 tokens.
@@
  • A-Mem [35] instantiates a memory system through multiple agents with dedicated prompts, including the memory retrieval construction, precursor memory evolution, and retrieval agents;
  • CAM [15] builds a hierarchical memory system via multiple agents, where dedicated agents undertake node construction (via ego-centric disentanglement and node replication) and information aggregation (for hierarchical abstraction). We implement it on the PersonaMem benchmark.
  • MemoryBank [42] construct a human-like memory for LLMs that uses a forgetting curve to remember and forget information over time. Similarly, we implement it on the PersonaMem benchmark.
  • SeCom [21] designs a modular memory system for long-horizon conversational agents, where dedicated components handle memory construction and retrieval using prompt-based multi-agent.
  • Mem0 [6] instantiates a multi-agent memory system that maintains fine-grained and coarse-grained user memories, performing extraction and retrieval with specialized prompts.
@@
  • GAM [37] introduces an agentic memory system with a memorizer for key information and a researcher for retrieval and integration at runtime, trained with reinforcement learning.
  • Mem1 [44] proposes a reinforcement learning (RL) framework that empowers agents to conduct efficient multi-turn information retrieval and reasoning for query answering. The optimized agents prioritize direct memory retrieval from history, rather than memory construction.
  • Memory-R1 [38] constructs a Memory Manager that learns optimal memory operations (ADD, UPDATE, DELETE) and an Answer Agent tasked with distilling relevant information from retrieved entries. Uniquely, the system is optimized via independent reinforcement learning (RL) in two stages: the Memory Manager is optimized first, followed by the Answer Agent.
@@ B.3 More Implementation Details
For a fair comparison, all the baselines are implemented with the same LLM backbones as our CoMAM, including Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct. All experiments are conducted on a cluster of 8 NVIDIA A100 GPUs. For the implementation of variant “+independent RL”, each agent is trained by fusing its task-specific rewards and global accuracy rewards, with the remaining agents frozen. For the “single policy” variant, we adopt a shared policy to serve as the policy for all three agents during training and inference. The prompts for specializing different agents and rubrics for evaluating the profile agent are presented below. The temperature of LLMs is set to $1$ for training and $0.8$ for evaluation. The detailed hyperparameter settings for CoMAM are presented in Table 8.
@@
Rubrics for Profile Agent Reward Evaluation You are an evaluation expert. Your task is to assess whether the provided Coarse-grained memory accurately captures the key information from the given Conversation History. Conversation History:
  User--LLM conversational history
Coarse-grained memory:
  Output generated by the Profile Agent
Evaluation Criteria: 1. Are the scenarios sufficiently rich and comprehensive? 2. Is the analysis of user preferences sufficiently in-depth? 3. Is the user persona sufficiently detailed and representative? Output Format: Please output a single scalar score within <score></score> as a float between 0.0 and 1.0, without any additional text. 1.0: The synthesized insight perfectly captures the scenario, user preferences, and persona. 0.0: The synthesized insight fails to capture the relevant information.
@@ C.1 The Impact of Joint Optimization on Local Task Specialization
  • Global performance enhancement: The global reward curve of “Ours” converges to 0.70, significantly higher than “w/ L’ (0.65) and “w/ G” (0.64), proving that joint optimization realizes global performance enhancement by fostering inter-agent collaboration.
Adaptive credit assignment mechanism quantifies the correlation between each agent’s local performance and the global outcome, ensuring that agents receive rewards not only for completing local tasks well but also for contributing to the global goal. This guides agents to optimize their local strategies in a way that benefits the entire system, thus achieving both local specialization and inter-agent collaboration for the global memory system.
C.2 Detailed Performance Across Question Types
Figure 6 details Qwen-based results for the seven question types, and we observe that CoMAM performs well in both detailed factual recall (e.g., Type 1) and abstract preference-based question types (e.g., Type 6), providing additional validation of its efficacy.
@@ D.1 Difference with existing memory framework
    Existing frameworks primarily adopt independent optimization. For example, MIRIX [29] designs a meta-agent and six specialized agents for memory management, optimizing each via tailored prompts; Memagent [39] focuses on a single retrieval agent, training it with query-answer accuracy as the reward; Memory-R1 [38] splits the system into a Memory Manager and Answer Agent, training them in separate stages—first optimizing the Manager with the Agent frozen, then freezing the Manager to optimize the Agent. This approach ignores inter-agent independencies and lacks co-adaptation: memory construction is not adjusted based on together with retrieval behavior, and the agent can only adapt to a fixed Manager’s output, leading to insufficient inter-agent collaboration.
    CoMAM proposes end-to-end joint optimization via modeling inter-agent dependencies in the MDP trajectory, where each agent’s output directly serves as the next agent’s input. All three agents are updated simultaneously during training, with inter-agent dependencies embedded into state transitions, ensuring they evolve collaboratively toward the global goal of the memory system.
@@
    Existing frameworks that optimize memory agents with RL primarily rely solely on final query-answer accuracy as the reward for all agents. For example, Mem-$\alpha$ and Mem1 [44] use this single reward to train construction and retrieval agents, respectively. Memory-R1 [38], and GAM [37] also adopt this design, using only answer accuracy to optimize their agents separately. This ambiguous credit assignment prevents agents from distinguishing their individual contributions, leading to unstable convergence.
@@ D.2 Difference with existing multi-agent system RL
  • Heterogeneous Policy vs. Homogeneous Policy: Some multi-agent reinforcement learning (RL) methods adopt a shared policy across all agents. For instance, CORY [18] proposes an RL framework featuring “Pioneer” and “Observer” agents that share a single policy model. These agents co-evolve through multi-agent collaboration and periodic role-swapping. In contrast, considering the heterogeneity of agents in a memory system, CoMAM employs heterogeneous policies tailored to distinct agents.
  • Workflow: Sequential vs. Concurrent: Existing multi-agent system (MAS) methods often adopt a debate-style pipeline for reasoning tasks, such as math tasks. For instance, MAPoRL [22] models agent interactions as concurrent debate, where multiple agents generate responses in parallel, focusing on persuasive collaboration for general reasoning. In contrast, CoMAM explicitly models inter-agent dependencies as a sequential Markov Decision Process (MDP), providing a structured way to expose how upstream decisions influence downstream behavior. This ensures that memory construction quality dynamically adapts to retrieval requirements and vice versa.
  • Collaboration Paradigm: Joint Optimization vs. Independent Alignment: Existing methods, such as Optimas [33], foster collaboration by aligning local rewards with global objectives via heuristic reward shaping for compound AI systems. However, it optimizes individual agents separately, leveraging locally-global aligned rewards rather than through end-to-end joint training. In contrast, CoMAM adopts a joint optimization paradigm that updates all agents simultaneously.
  • Credit Assignment: Adaptive Assignment vs. Equal Allocation: Most existing systems use static credit assignment mechanisms without quantifying agents’ varying contributions: AT-GRPO [40] distributes rewards equally across agents using fixed weights, ignoring task differences; CoMAS [36] adopts a zero-sum adversarial reward scheme that pits agents against each other, which is misaligned with the cooperative nature of memory systems; In contrast, CoMAM assigns agent-specific credit based on a ranking-based proxy that measures the alignment between each agent’s local task rewards and the global system reward. This formulation captures whether trajectories with higher global performance are consistently associated with higher local rewards for a given agent. Agents with stronger alignment receive higher credit, as their behaviors are more predictive of overall system improvement. This provides more informative training signals for specialized agents.
@@ Appendix E Limitation
While CoMAM demonstrates the effectiveness of joint optimization with adaptive credit assignment for multi-agent memory systems, several limitations remain.
1. Simplified Agent Design. We adopt a compact set of agents (Extraction, Profile, and Retrieval) to focus on core collaborative behaviors. More advanced functionalities, such as memory editing, consistency maintenance, and redundancy removal, are not explicitly modeled and remain for future exploration.
2. Limited Cross-Agent Credit Modeling. CoMAM estimates agent-specific credit by modeling the ranking consistency between each agent’s local rewards and the global system reward. However, relationships among agents’ local rewards are not explicitly considered in the credit estimation. Incorporating such cross-agent relationships may provide more informative credit signals and further improve coordination.
Appendix F Broader Impact
CoMAM aims to improve the effectiveness of multi-agent memory systems for long-horizon conversational tasks by introducing a joint optimization framework with end-to-end reinforcement learning and adaptive credit assignment. This may benefit applications such as personalized assistants and long-context reasoning systems. However, it does not inherently address potential privacy risks associated with sensitive user information.